OpenStudy (anonymous):

What is the diameter of a circle whose circumference is 89.49 centimeters? Use 3.14 for π.

jimthompson5910 (jim_thompson5910):

C = pi*d C = circumference d = diameter

jimthompson5910 (jim_thompson5910):

C = pi*d 89.49 = 3.14*d solve for d and tell me what you get